# version 0.75

### v4

@@ 4.3 Controllability of LDS
The controllability of a linear dynamical system can now be characterized using this matrix. Recall the rank of a matrix is the maximal number of linearly independent columns (or rows).
@@
Consider the evolution of the noiseless linear dynamical system starting from initial state $\mathbf{x}_{0}$. The state at time $T$ is given by the expression
@@
Adding ${B}$ to both matrices doesn’t affect the subset relationship.
@@
Therefore, $r_{t+2}\leq r_{t+1}$. Since the rank cannot decrease through the addition of columns, we have $r_{t+2}=r_{t+1}$. Since the rank can grow at most $d_{x}$ times, the equation $r_{d_{x}}=r_{T}$ holds for all $T>d_{x}$ if and only if the system is controllable, and this concludes the theorem. ∎
@@ 4.4 Quantitative Definitions
In the rest of this text our goal is to give efficient algorithms with finite time performance bounds for online control. These finite time bounds naturally depend on how difficult it is for a given system to stabilize. Therefore, we require a quantitative definition for stabilizability, one that holds for both time-invariant and time-varying linear dynamical systems.
4.4.1 Stabilizability of Time-Invariant LDS
Consider a linear time-invariant system given by the transformations $({A,B})$. We have already remarked that this system is called stabilizable, according to the classical, noise-free definition, if and only if there exists $K$ such that the condition $\rho({A+BK})<1$ holds.
A stronger quantitative property can be obtained as follows, and it is in fact implied from BIBO stabilizability as defined in Definition 2.4. The proof of this fact is left as exercise 4.6.
@@ 4.4.1 Stabilizability of Time-Invariant LDS
Consider a linear controller ${K}$ for the LTI system $({A,B})$. If there exists $\kappa>0$ and $\delta\in(0,1]$ such that the matrix ${A+BK}$ admits the following inequality for all $t$, then $K$ is a $\gamma$-stabilizing controller for $(A,B)$.
@@
A linear time-invariant system is thus $\gamma$-stabilizable if we can find a matrix $K$ with the properties above.
4.4.2 Stabilizability of time-variant linear systems
@@
We now define a sufficient condition for $\gamma$ stabilizability. Consider a time-varying linear dynamical system given by the transformations $\{{A}_{t},{B}_{t}\}$.
@@ 4.4.3 Controllability of time-invariant systems
Clearly, such a system is nominally controllable, as indeed is any one-dimensional system with non-zero $B$. However, to reach a state $x=1$ from an initial $x_{0}=0$ in a constant number of time steps requires a large control input $u_{t}$ of size at least $1/\varepsilon$. Thus, such a system may be said to be barely controllable. This difficulty of controllability is captured in the definition below, which requires not only that $K(A,B)$ be full row-rank, but also that such a statement is robust to small changes in $A,B$.
@@
$\mathop{\mbox{\rm rank}}{K}_{k}({A},{B})=d_{x},\text{ and }\|({K}_{k}(A,B))^{-1}\|\leq\kappa.$
In the definition above, the spectral norm of the inverse of $K(A,B)$ is bounded by $\kappa$. For example, in the simple one-dimensional system we presented above, the matrix $K(A,B)$ is $\varepsilon$, and thus $\kappa=\frac{1}{\varepsilon}$, which indeed captures the complexity of controlling the system.
@@ 4.5 Bibliographic Remarks
The idea of reducing a nonlinear dynamical system to a time varying linear dynamical system had been independently discovered in disparate fields. In the context of control, using the linearization to characterize local stability properties of non-linear systems was pioneered by Lyapunov [Lya92], and is called the Lyapunov direct method, see e.g. [SL91].
@@
The efficient verification of stabilizability via a semi-definite program and the recovery of a stable controller form the solution of the latter appears in [Ahm16]. On the other hand, it is known that even imposition of norm constraints on a plausibly stable controller makes this problem computationally hard [BT97, BT00, BT99]. Somewhat distinct, but related, problems of stabilizability with even a 1-dimensional linear constraint are open to the date.
The notion of $\gamma$-stabilizability has a precursor in the notion of strong stability introduced in [CHK+18]. An extension of this strong stability notion to time-varying linear dynamics was established in [CKM19]. The definition presented in this chapter has weaker requirements than those in the listed articles, allowing us to work with and guarantee regret later with respect to a larger class of comparator policies.
@@ 4.6 Exercises
4.1. Prove that the existence of a matrix $K$ such that $\rho(A+BK)<1$ is a necessary condition for a linear dynamical system to be stabilizable. Hint: this exercise requires the optimal solution to LQR from the next chapter.
@@ 4.4. Prove that a stochastic linear dynamical system with Gaussian perturbations
Hint: One path to this exercise is to explicitly calculate the stationary distribution of a LDS and then to verify reversibility as stated in its definition. Exercise 3.6 provides an alternative without the need to derive the stationary distribution.
4.5. This exercise explores the consequences of $\gamma$-BIBO stability (Definition 2.4) for linear systems. While we have already seen sufficient conditions to ensure BIBO stability, we ask for the necessity of such conditions.
A. For a time-invariant linear system $(A,B)$ equipped with a linear policy $K$, prove that BIBO stability indeed implies that $\rho(A+BK)<1$. In fact, comment on how large $\rho(A+BK)$ can be.
B. For a time-invariant BIBO-stable linear system, prove that there exists $\kappa>0$ and $\delta\in(0,1]$ such that, for all $t$, $\|(A+BK)^{t}\|\leq\kappa(1-\delta)^{t}$. In particular, this proves that the conditions listed in Lemma 4.8 are necessary, in addition to being sufficient.
C. Construct a time-varying linear system that is BIBO stable, yet does not have the exponential decay properties required in Lemma 4.9.
4.6. Prove that the bound of one on the operator norm of the matrix $A$ is sufficient but not necessary condition for stabilizability.
@@ 5.1 The Linear Quadratic Regulator
To remind the reader, the notation we use is as follows:
@@
This framework is very useful since many physical problems, such as the centrifugal governor and the ventilator control problem, can be approximated by a linear dynamical system. In addition, the quadratic convex cost functions are natural in many physical applications.
The LQR is a special case of a Markov decision process which we have described in chapter 3, since the transitions are stochastic. However, we cannot naive apply value iteration, policy iteration, or linear programming since there are infinitely many states and actions.
@@ 5.2 Optimal Solution of the LQR
We know from Chapter 3 that the LQR is a special case of a finite-time Markov Decision Process, and hence admits an optimal policy. However, the optimal policy is in general a mapping from states to actions, both of which are infinite Euclidean spaces in our case.
Nevertheless, it can be shown that the optimal policy for an LQR is a linear policy, i.e. a linear mapping from ${\mathbb{R}}^{d_{x}}$ to ${\mathbb{R}}^{d_{u}}$! This astonishing fact, together with a remarkable property of the value function, is formally given in the following theorem. We state the theorem for a finite number of iterations, and for a linear time-invariant system. Generalizing this to a time-varying linear system is left as exercise 5.6.
@@
Since $R$ and $S_{t}$ are positive semi-definite, the expression above is convex as a function of $\mathbf{u}$. At the global minimum the gradient vanishes, and hence
@@ 5.3 Infinite Horizon LQR
Although infinite horizon problems can never appear in the real world, in practice the horizon length may be unknown. Although there are techniques to iteratively update the horizon length with geometric increasing guesstimates, these are slow and impractical.
@@
where similar to the finite horizon case, we assume that the perturbations $\mathbf{w}_{t}$ are i.i.d. zero mean and have bounded variance. We henceforth assume that the perturbations are all zero for simplicity, adding zero-mean bounded variance perturbations is left as exercise 5.6.
@@
Similar to the finite time case, we have the following result.
@@
At optimality, the gradient with respect to $\mathbf{u}$ of the equation above is zero. This implies that
@@
Plugging back into the Bellman equation, we obtain
@@ 5.3.1 Solving the DARE
In the 1-D case, although the right side of the equation is a rational function, we can multiply both sides by ${R+B^{\top}SB}$ to form a quadratic equation in ${S}$. Solving the DARE therefore corresponds to finding a non-negative root of a quadratic polynomial in the one-dimensional case. The reason we only consider the non-negative root is since the value function is non-negative by our choice of quadratic convex costs.
@@ 5.4 Robust Control
One of the most important and prominent generalizations is that of robust control, also known as $\text{H}_{\infty}$-control. The goal of $\text{H}_{\infty}$-control is to come up with a more robust solution, by considering the optimal solution vs. any possible perturbation sequence from a certain family. Formally, we aim to solve:
@@
Notice that this problem does not fall into the category of a Markov Decision Process, since $\mathbf{w}_{t}$’s are no longer stochastic. Therefore the typical Bellman optimality criterion does not apply.
Instead, the appropriate techniques to address this problem come from the theory of dynamical games. The above is, in fact, a two-player zero sum game. The first player tries to minimize the costs and the other attempts to maximize the costs, both players subject to the dynamics evolution constraint.
@@
The proof is somewhat more involved than the LQR derivation, and is omitted from this text.
@@ 5.5 Bibliographic Remarks
While seemingly a natural corollary of equivalent results for MDPs, the existence of value functions for linear control is made subtle by the continuous nature of the state space and stabilizability, or the lack thereof. A discussion on sufficient and necessary conditions to existence of solutions to Riccati equations can be found in [Ber07, HSK96].
The development of $H_{\infty}$ control was driven by an observation made in the seminal paper of Doyle [Doy78] that, unlike in the fully observed case, the solution to the optimal control problem for partially observed linear systems (which we study in later chapters) is not naturally robust to miss-specification of system parameters. This suggests that robustness specifications constitute yet an another axis of performance for control algorithms, and may at times be at odds with the objective of optimal control. For in-depth treatment of $H_{\infty}$ control, see [ZDG96b]. The text [BB08] presents a succinct derivation of robust controllers entirely in the state space, as the solution to a dynamic game between the controller and an adversary, see also [Tu99] for a recent exposition.
@@ 5.6 Exercises
5.2. Derive the solution to the infinite horizon LQR with zero mean Gaussian noise.
5.3. Extend the proof of Theorem 5.2 to positive semi-definite $Q,R$, rather than positive definite matrices.
5.4. Prove Theorem 5.4 as a dynamic two-player zero-sum game, where a unique saddle point solution exists. Hint: you can formulate this problem as a min-max optimization problem within a finite horizon.
5.5. Prove that the function $f(\mathbf{x})=\mathbf{x}^{\top}M\mathbf{x}$ is convex if and only if $M\succeq 0$ is positive semi-definite.
@@
In the first part of this book and chapter 5 we studied linear dynamical systems with quadratic costs, namely the LQR problem. We saw that in this case the optimal control takes a very special form: it is a linear function of the state. The Markovian property of the Markov Decision Process framework for reinforcement learning suggests that the optimal policy is always a function of the current state. It is characterized by the Bellman optimality equation, and can be computed in complexity proportional to the number of states and actions using, for example, the value iteration algorithm which we analyzed in a previous chapter.
@@
In such scenarios involving significant deviations from the MDP framework, classical notions of optimality may be ill-defined. Or if they do exist, the optimal policy may no longer be a mapping from immediate state to action. Rather, the learner may be required to use other signals available to her when choosing an action that is most beneficial in the long run. To this we add a significant consideration: computation of optimal state-based policies may be computationally intractable.
This motivates the current chapter which considers regret analysis as a computational tractable alternative to optimal or robust control.
We start by considering regret in the decision making framework of online convex optimization. We then consider several classes of control policies, as well as the relationships between them, in the context of regret analysis.
@@ 6.1 Online Convex Optimization
In this section we give a brief introduction to the theory of Online Convex Optimization. The reader is referred to the bibliographic section for a comprehensive survey.
In the decision making framework of Online Convex Optimization, a player iteratively chooses a point in a convex set $\mathcal{K}\subseteq{\mathbb{R}}^{n}$, and suffers a loss according to an adversarially chosen loss function $f_{t}:\mathcal{K}\mapsto{\mathbb{R}}$.
@@
If the algorithm is clear from the context, we henceforth omit the superscript and denote the algorithm’s decision at time $t$ simply as $\mathbf{x}_{t}$. Intuitively, an algorithm performs well if its regret is sublinear as a function of $T$ (i.e. $\mathrm{{regret}}_{T}({\mathcal{A}})=o(T)$), since this implies that on average, the algorithm performs as well as the best fixed strategy in hindsight. We proceed to describe a basic algorithm for online convex optimization.

### v5

@@ 4.3 Controllability of Linear Dynamical Systems
The controllability of a linear dynamical system can now be characterized using this matrix. Recall that the rank of a matrix is the maximal number of linearly independent columns (or rows).
@@
Consider the evolution of the noiseless linear dynamical system starting from the initial state $\mathbf{x}_{0}$. The state at time $T$ is given by the expression
@@
Adding ${B}$ to both matrices does not affect the subset relationship.
@@
Therefore, $r_{t+2}\leq r_{t+1}$. Since the rank cannot decrease through the addition of columns, we have $r_{t+2}=r_{t+1}$. Since the rank can grow at most $d_{x}$ times, equation $r_{d_{x}}=r_{T}$ holds for all $T>d_{x}$ if and only if the system is controllable, and this concludes the theorem. ∎
@@ 4.4 Quantitative Definitions
In the rest of this text our goal is to give efficient algorithms with finite time performance bounds for online control. These finite-time bounds naturally depend on how difficult it is for a given system to stabilize. Therefore, we require a quantitative definition for stabilizability, one that holds for both time-invariant and time-varying linear dynamical systems.
4.4.1 Stabilizability of Time-Invariant Systems
Consider a linear time-invariant system given by the transformations $({A,B})$. We have already remarked that this system is called stabilizable, according to the classical definition, if and only if there exists $K$ such that $\rho({A+BK})<1$ holds.
A stronger quantitative property can be obtained as follows, and in fact it is implied by BIBO stabilizability as defined in Definition 2.4. The proof of this fact is left as exercise 4.6.
@@ 4.4.1 Stabilizability of Time-Invariant Systems
Consider a linear controller ${K}$ for the LTI system $({A,B})$. If there exist $\kappa>0$ and $\delta\in(0,1]$ such that the matrix ${A+BK}$ admits the following inequality for all $t$, then $K$ is a $\gamma$-stabilizing controller for $(A,B)$.
@@
A linear time-invariant system is thus $\gamma$-stabilizable if we can find a matrix $K$ with the above properties.
4.4.2 Stabilizability of Time-Varying Systems
@@
We now define a sufficient condition for $\gamma$-stabilizability. Consider a time-varying linear dynamical system given by the transformations $\{{A}_{t},{B}_{t}\}$.
@@ 4.4.3 Controllability of Time-Invariant Systems
Clearly, such a system is nominally controllable, as is indeed any one-dimensional system with nonzero $B$. However, reaching a state $x=1$ from an initial $x_{0}=0$ in a constant number of time steps requires a large control input $u_{t}$ of magnitude at least $1/\varepsilon$. Thus, such a system may be said to be barely controllable. This difficulty of controllability is captured in the definition below, which requires not only that $K(A,B)$ is a full row-rank, but also that such a statement is robust to small changes in $A,B$.
@@
$\mathop{\mbox{\rm rank}}{K}_{k}({A},{B})=d_{x},\text{ and }\|({K}_{k}(A,B){K}_{k}(A,B)^{\top})^{-1}\|\leq\kappa.$
In the above definition, in addition to the full row-rank constraint, the spectral norm of the inverse of the Gram matrix of $K(A,B)$ is bounded by $\kappa$. For example, in the simple one-dimensional system we presented above, with $k=1$, the matrix $K(A,B)$ is $\varepsilon$, and thus $\kappa=\frac{1}{\varepsilon^{2}}$, which indeed captures the complexity of controlling the system.
@@ 4.5 Bibliographic Remarks
The idea of reducing a nonlinear dynamical system to a time varying linear dynamical system had been independently discovered in disparate fields. In the context of control, using the linearization to characterize local stability properties of non-linear systems was pioneered by Lyapunov [Lya92], and is called the Lyapunov direct method; see, e.g. [SL91].
@@
The efficient verification of stabilizability via a semidefinite program and the recovery of a stable controller form the solution of the latter appears in [Ahm16]. On the other hand, it is known that even imposition of norm constraints on a plausibly stable controller makes this problem computationally hard [BT97, BT00, BT99]. Somewhat distinct but related problems of stabilizability with even a 1-dimensional linear constraint are open to the date.
The notion of $\gamma$-stabilizability has a precursor in the notion of strong stability introduced in [CHK+18]. An extension of this strong stability concept to time-varying linear dynamics was established in [CKM19]. The definition presented in this chapter has weaker requirements than those in the listed articles, allowing us to work with and guarantee regret later with respect to a larger class of comparator policies.
@@ 4.6 Exercises
4.1. Prove that the existence of a matrix $K$ such that $\rho(A+BK)<1$ is a necessary condition for a linear dynamical system to be stabilizable. Hint: this exercise requires the optimal solution to LQR in the next chapter.
@@ 4.4. Prove that a stochastic linear dynamical system with Gaussian perturbations
Hint: One way to do this exercise is to explicitly calculate the stationary distribution of an LDS and then to verify reversibility as stated in its definition. Exercise 3.6 provides an alternative without the need to derive the stationary distribution.
4.5. This exercise explores the consequences of $\gamma$-BIBO stability (Definition 2.4) for linear systems. Although we have already seen sufficient conditions to ensure BIBO stability, we ask for the necessity of such conditions.
A. For a time-invariant linear system $(A,B)$ equipped with a linear policy $K$, prove that BIBO stability indeed implies $\rho(A+BK)<1$. In fact, comment on how large $\rho(A+BK)$ can be.
B. For a time-invariant BIBO-stable linear system, prove that there exist $\kappa>0$ and $\delta\in(0,1]$ such that, for all $t$, $\|(A+BK)^{t}\|\leq\kappa(1-\delta)^{t}$. In particular, this proves that the conditions listed in Lemma 4.8 are necessary, in addition to being sufficient.
C. Construct a time-varying linear system that is BIBO stable but does not have the exponential decay properties required in Lemma 4.9.
4.6. Prove that the bound of one on the operator norm of the matrix $A$ is a sufficient but not necessary condition for stabilizability.
@@ 5.1 The Linear Quadratic Regulator
To remind the reader, the notation we use is as follows.
@@
This framework is very useful, since many physical problems, such as the centrifugal governor and the ventilator control problem, can be approximated by a linear dynamical system. In addition, quadratic convex cost functions are natural in many physical applications.
The LQR is a special case of a Markov decision process which we have described in Chapter 3, since the transitions are stochastic. However, we cannot naively apply value iteration, policy iteration, or linear programming since there are infinitely many states and actions.
@@ 5.2 Optimal Solution of the LQR
We know from Chapter 3 that the LQR is a special case of a finite-time Markov decision process and hence admits an optimal policy. However, the optimal policy is in general a mapping from states to actions, both of which are infinite Euclidean spaces in our case.
However, it can be shown that the optimal policy for an LQR is a linear policy, that is, a linear mapping from ${\mathbb{R}}^{d_{x}}$ to ${\mathbb{R}}^{d_{u}}$! This astonishing fact, together with a remarkable property of the value function, is formally given in the following theorem. We state the theorem for a finite number of iterations and for a linear time-invariant system. Generalizing this to a time-varying linear system is left as exercise 5.6.
@@
Since $R$ and $S_{t}$ are positive semidefinite, the above expression is convex as a function of $\mathbf{u}$. At the global minimum the gradient vanishes, and hence
@@ 5.3 Infinite Horizon LQR
Although infinite horizon problems can never appear in the real world, in practice the horizon length may be unknown. Although there are techniques to iteratively update the horizon length with geometrically increasing guesstimates, these are slow and impractical.
@@
where similar to the finite horizon case, we assume that the perturbations $\mathbf{w}_{t}$ are i.i.d. zero mean and have bounded variance. We henceforth assume that the perturbations are all zero for simplicity, and adding zero-mean bounded-variance perturbations is left as exercise 5.6.
@@
Similarly to the finite-time case, we have the following result.
@@
At optimality, the gradient with respect to $\mathbf{u}$ of the above equation is zero. This implies that
@@
Plugging this back into the Bellman equation, we obtain
@@ 5.3.1 Solving the DARE
In the one-dimensional case, although the right side of the equation is a rational function, we can multiply both sides by ${R+B^{\top}SB}$ to form a quadratic equation in ${S}$. Solving the DARE therefore corresponds to finding a nonnegative root of a quadratic polynomial in the one-dimensional case. The reason we only consider the nonnegative root is that the value function is nonnegative by our choice of quadratic convex costs.
@@ 5.4 Robust Control
One of the most important and prominent generalizations is that of robust control, also known as $\text{H}_{\infty}$-control. The goal of $\text{H}_{\infty}$-control is to come up with a more robust solution, by considering the optimal solution vs. any possible perturbation sequence from a certain family. Formally, our aim is to solve the following:
@@
Notice that this problem does not fall into the category of a Markov Decision Process, since $\mathbf{w}_{t}$’s are no longer stochastic. Therefore, the typical Bellman optimality criterion does not apply.
Instead, the appropriate techniques to address this problem come from the theory of dynamical games. The above is, in fact, a two-player zero sum game. The first player tries to minimize the costs, and the other attempts to maximize the costs, both players subject to the dynamics evolution constraint.
@@
The proof is somewhat more involved than the LQR derivation and is omitted from this text.
@@ 5.5 Bibliographic Remarks
While seemingly a natural corollary of equivalent results for MDPs, the existence of value functions for linear control is made subtle by the continuous nature of the state space and stabilizability, or the lack thereof. A discussion of sufficient and necessary conditions for the existence of solutions to Riccati equations can be found in [Ber07, HSK96].
The development of $H_{\infty}$ control was driven by an observation made in the seminal paper of Doyle [Doy78] that, unlike in the fully observed case, the solution to the optimal control problem for partially observed linear systems (which we study in later chapters) is not naturally robust to misspecification of system parameters. This suggests that robustness specifications constitute yet another axis of performance for control algorithms and may at times be at odds with the objective of optimal control. For in-depth treatment of $H_{\infty}$ control, see [ZDG96b]. The text [BB08] presents a succinct derivation of robust controllers entirely in the state space, as the solution to a dynamic game between the controller and an adversary; see also [Tu18] for a recent exposition.
@@ 5.6 Exercises
5.2. Derive the solution to the infinite-horizon LQR with zero mean Gaussian noise.
5.3. Extend the proof of Theorem 5.2 to positive semidefinite $Q,R$, rather than positive definite matrices.
5.4. Prove Theorem 5.4 as a dynamic two-player zero-sum game, where a unique saddle-point solution exists. Hint: you can formulate this problem as a min-max optimization problem within a finite horizon.
5.5. Prove that the function $f(\mathbf{x})=\mathbf{x}^{\top}M\mathbf{x}$ is convex if and only if $M\succeq 0$ is positive semidefinite.
@@
In the first part of this book and in chapter 5 we studied linear dynamical systems with quadratic costs, namely the LQR problem. We saw that in this case the optimal control takes a very special form: it is a linear function of the state. The Markovian property of the Markov Decision Process framework for reinforcement learning suggests that the optimal policy is always a function of the current state. It is characterized by the Bellman optimality equation, and can be computed in complexity proportional to the number of states and actions using, for example, the value iteration algorithm, which we analyzed in a previous chapter.
@@
In such scenarios involving significant deviations from the MDP framework, classical notions of optimality may be ill-defined. Or if they do exist, the optimal policy may no longer be a mapping from immediate state to action. Rather, the learner may be required to use other signals available to her when choosing an action that is most beneficial in the long run. To this we add a significant consideration: the computation of optimal state-based policies may be computationally intractable.
This motivates the current chapter, which considers regret analysis as a computationally tractable alternative to optimal or robust control.
We start by considering regret in the decision-making framework of online convex optimization. We then consider several classes of control policies and the relationships between them in the context of regret analysis.
@@ 6.1 Online Convex Optimization
In this section, we give a brief introduction to the theory of Online Convex Optimization. The reader is referred to the bibliographic section for a comprehensive survey.
In the decision-making framework of Online Convex Optimization, a player iteratively chooses a point in a convex set $\mathcal{K}\subseteq{\mathbb{R}}^{n}$, and suffers a loss according to an adversarially chosen loss function $f_{t}:\mathcal{K}\mapsto{\mathbb{R}}$.
@@
If the algorithm is clear from the context, we henceforth omit the superscript and denote the algorithm’s decision at time $t$ simply as $\mathbf{x}_{t}$. Intuitively, an algorithm performs well if its regret is sublinear as a function of $T$ (i.e. $\mathrm{{regret}}_{T}({\mathcal{A}})=o(T)$), since this implies that, on average, the algorithm performs as well as the best fixed strategy in hindsight. We then describe a basic algorithm for online convex optimization.